# Strongly coupled gauge theories towards physics beyond the Standard Model

### v1

@@ 2.1 Conformal window
Renormalization group (RG) equations describe the evolution of a physical system with respect to the energy scale, and have been used to understand how long-distance or infra-red (IR) dynamics emerge from short-distance or ultra-violet (UV) dynamics. In particular, the RG beta function $\beta(\alpha)=\frac{\partial\alpha}{\partial\ln\mu}$ of quantum field theory encodes the change of the coupling constant $\alpha$ with the renormalized scale $\mu$. If we restrict ourselves to non-abelian gauge theories coupled to $N_{f}$ massless fermions with a coupling constant $g$, the perturbative calculation at one-loop immediately finds $\beta(g)<0$ for $N_{f}<N_{f}^{\rm AF}$ and thus yields asymptotic freedom with a gaussian fixed point at $g=0$. While theories with sufficiently small $N_{f}$, including Yang-Mills, are supposed to be confined and chirally broken, theories with $N_{f}$ just below $N_{f}^{\rm AF}$ (assuming fractional numbers of $N_{f}$) develop a perturbative IR fixed point at two-loop order, called Banks-Zaks (BZ) fixed point. Without loss of generality, asymptotic free gauge theories are therefore further divided into two distinct phases in which the IR features are completely different. A conformal window (CW) is defined as the range of $N_{f}$ over which the theories are IR conformal, i.e. $N_{f}^{\rm cr}<N_{f}<N_{f}^{\rm AF}$.
@@
The perturbative analysis of the conformal window can be used to make an estimate of $N_{f}^{\rm cr}$ by imposing that the coupling at the BZ fixed point runs to infinity, which can be improved by adding higher loop-order corrections to the beta function, $\beta(\alpha)=-2\alpha\sum_{\ell=1}^{\infty}b_{\ell}\left(\frac{\alpha}{4\pi}\right)^{\ell}$. However, this approach becomes not only scheme dependent for $\ell\geq 3$, but also results in a severe convergence problem when the $5$th order results in the $\overline{\rm MS}$ scheme, the highest loop-order known to us so far, are concerned [6]. An alternative scheme-independent expansion in $\Delta_{N_{f}}=N_{f}^{\rm AF}-N_{f}^{\rm IR}$, called conformal or BZ expansion, has been proposed [7] and, for the past few years, extensively used to compute certain physical observables of IR conformal gauge theories by Ryttov and Shrock [8, 9, 10, 11, 12, 13, 14, 15]. A particularly interesting quantity is the anomalous dimension of a fermion condensate
@@ 5 Other topics
Another aspect of the SYM is that a certain sector of the theory is connected to its counterpart of non-supersymmetric field theories through the large $N$ equivalence. For instance, it has been argued that the bosonic sector of a gauge theory coupled to a Dirac fermion in the antisymmetric representation is non-perturbatively equivalent to that of $\mathcal{N}=1$ SYM in the large $N$ limit (orientifold equivalence) [106]. This has motivated recent lattice studies on $SU(N)$ gauge theories with one flavor of antisymmetric Dirac fermions. The first result with $N=3$ was reported in Ref. [107], where the meson spectrum has been computed and compared with the analytical predictions derived using an effective field theory approach. Extensions to a larger $N$ are currently underway: some preliminary results and technical challenges were discussed at this conference [108].
@@ Acknowledgments
I would like to thank S. Catterall, G. Catumba, A. Hasenfratz, C.-J. D. Lin, B. Lucini, A. Lupo, D. Mason, M. D. Morte, D. Nogradi, C. Peterson, D. Schaich, Y. Shamir and R. Shrock for providing material in advance of my talk and for fruitful discussions. J. W. L. acknowledges support by IBS under the project code, IBS-R018-D1.

### v2

@@ 2.1 Conformal window
Renormalization group (RG) equations describe the evolution of a physical system with respect to the energy scale, and have been used to understand how long-distance or infra-red (IR) dynamics emerge from short-distance or ultra-violet (UV) dynamics. In particular, the RG beta function $\beta(\alpha)=\frac{\partial\alpha}{\partial\ln\mu}$ of quantum field theory encodes the change of the coupling constant $\alpha$ with the renormalized scale $\mu$. If we restrict ourselves to non-abelian gauge theories coupled to $N_{f}$ massless fermions with a coupling constant $g$, the perturbative calculation at one-loop immediately finds $\beta(g)<0$ for $N_{f}<N_{f}^{\rm AF}$ and thus yields asymptotic freedom with a gaussian fixed point at $g=0$ [4, 5]. While theories with sufficiently small $N_{f}$, including Yang-Mills, are supposed to be confined and chirally broken, theories with $N_{f}$ just below $N_{f}^{\rm AF}$ (assuming fractional numbers of $N_{f}$) develop a perturbative IR fixed point at two-loop order [6], called Banks-Zaks (BZ) fixed point. Without loss of generality, asymptotic free gauge theories are therefore further divided into two distinct phases in which the IR features are completely different. A conformal window (CW) is defined as the range of $N_{f}$ over which the theories are IR conformal, i.e. $N_{f}^{\rm cr}<N_{f}<N_{f}^{\rm AF}$.
@@
The perturbative analysis of the conformal window can be used to make an estimate of $N_{f}^{\rm cr}$ by imposing that the coupling at the BZ fixed point runs to infinity, which can be improved by adding higher loop-order corrections to the beta function, $\beta(\alpha)=-2\alpha\sum_{\ell=1}^{\infty}b_{\ell}\left(\frac{\alpha}{4\pi}\right)^{\ell}$, e.g. see Ref. [9]. However, this approach not only becomes scheme-dependent for $\ell\geq 3$, but also leads to serious convergence problems when the $5$th order results in the $\overline{\rm MS}$ scheme, the highest loop-order known to us so far, are concerned [10]. An alternative scheme-independent expansion in $\Delta_{N_{f}}=N_{f}^{\rm AF}-N_{f}^{\rm IR}$, called conformal or BZ expansion, has been considered in the context of (asymptotically free) IR conformal gauge theories [11, 12] and, for the past few years, extensively used to compute certain physical observables at an IR fixed point by Ryttov and Shrock [13, 14, 15, 16, 17, 18, 19, 20]. A particularly interesting quantity is the anomalous dimension of a fermion condensate
@@ 5 Other topics
Another aspect of the SYM is that a certain sector of the theory is connected to its counterpart of non-supersymmetric field theories through the large $N$ equivalence. For instance, it has been argued that the bosonic sector of a gauge theory coupled to a Dirac fermion in the antisymmetric representation is non-perturbatively equivalent to that of $\mathcal{N}=1$ SYM in the large $N$ limit (orientifold equivalence) [111]. Together with a number of predictions made in Ref. [112], this has motivated recent lattice studies of $SU(N)$ gauge theory with one flavor of antisymmetric Dirac fermions. The first result with $N=3$ was reported in Ref. [113], where the meson spectrum has been computed and compared with the analytical predictions derived using an effective field theory approach. Extensions to a larger $N$ are currently underway: some preliminary results and technical challenges were discussed at this conference [DellaMorte:2023sdz].
@@ Acknowledgments
I would like to thank S. Catterall, G. Catumba, A. Hasenfratz, C.-J. D. Lin, B. Lucini, A. Lupo, D. Mason, M. D. Morte, D. Nogradi, C. Peterson, D. Schaich, Y. Shamir and R. Shrock for providing material in advance of my talk and for fruitful discussions. I also thank F. Sannino for helpful comments and suggestions on the manuscript. J. W. L. acknowledges support by IBS under the project code, IBS-R018-D1.